OpenStudy (anonymous):

Find the shortest distance from the origin to the parabola y=4-x^2

OpenStudy (anonymous):

let (x,y) be the point that yields the minimum distance from the origin

OpenStudy (anonymous):

\[D^2=(x-x_0)^2+(y-y_0)^2\] \[(x_0,y_0) \implies (0,0)\] \[D^2=(x-0)^2+(y-0)^2=x^2+y^2\]

OpenStudy (anonymous):

since \[y=4-x^2\] then \[y^2=(4-x^2)^2=16-8x^2+x^4\] so\[D^2(x)=x^4-8x^2+16+x^2=x^4-7x^2+16\] minimise \[D^2(x)\] by finding the derivative